OpenStudy (anonymous):

which set of values can be the side length of a 30 60 90 triangle?

OpenStudy (anonymous):

I'm not sure of your question: the values of the sides have to follow the formula of 30-60-90 triangles. I'm not sure if you want me to help you chose an answer if there are no options

OpenStudy (anonymous):

Short leg = (1/2)hypotenuse Long leg = (1/2)hypotenuse *sqrt(3)

OpenStudy (anonymous):

A.{4,4r2,8} B.{4,4r3,8} C.{4,4r3,8r3} D.{4,4r2,8r2 sorry it took me so long this is my question here and i wasn't sure how to put radical in the equation so i just put r

OpenStudy (anonymous):

I'm guessing it goes in the order (short leg, long leg, hypotenuse) ?

OpenStudy (anonymous):

yeah its on a site called apex learning

OpenStudy (anonymous):

radical is what i meant

OpenStudy (anonymous):

so 4r2 = \[4\sqrt{3}\]

OpenStudy (anonymous):

*4r3

OpenStudy (anonymous):

how do u enter the radical symbol in the equation

OpenStudy (anonymous):

There's a blue button that says "equation" under the chat bar.

OpenStudy (anonymous):

Use the forumla I posted above

OpenStudy (anonymous):

Short leg = 0.5(hypotenuse) Long leg = 0.5(hypotenuse)*\[\sqrt{3}\]

OpenStudy (anonymous):

also, the hypotenuse is the longest side :P

OpenStudy (anonymous):

is there a way i could just send a picture of my screen... it would be a lot easier for you

OpenStudy (anonymous):

I know what you mean and the answer, don't worry about me.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you have your answer?

OpenStudy (anonymous):

no not yet

OpenStudy (anonymous):

Did you try to use what I posted above? Once you do it's very easy

OpenStudy (anonymous):

sorry but im not sure how to use it

OpenStudy (anonymous):

You know the short leg is 4, because that's the answer for all the options

OpenStudy (anonymous):

oh your right

OpenStudy (anonymous):

Find hypotenuse by multiplying by 2, and find the long leg by multiplying by sqrt(3)

OpenStudy (anonymous):

i dont know which number to multiply by 2 it doesn't show the hypotenuse

OpenStudy (anonymous):

The order goes like this: shortleg, shortleg*sqrt(3), shortleg*2

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

do you have the answer now :P

OpenStudy (anonymous):

workin on it...im just confused

OpenStudy (anonymous):

Here's a hint, because square root 2 isn't in the formula, you can eliminate all answers with square root 2

OpenStudy (anonymous):

soo that just leaves B and C right ?

OpenStudy (anonymous):

Yes, and short leg is 4, because 4 is the first number for all the options

OpenStudy (anonymous):

and hypotenuse = short leg*2, so the last number should not have a radical

OpenStudy (anonymous):

took a while, but I think you know the answer now :)

OpenStudy (anonymous):

oh please tell me B is the right answer

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

@$#!@ yeahh!!! thnx for your help